### v1

# Thermal condensation with polybaric pressure

Abstract
In this work, we have carried out a two-dimensional (2D) simulation of thermal instability (TI) in interstellar matter (ISM), considering it to be a weakly ionized inviscid plasma with radiation loss. We carry out the simulation with the help of our multi-fluid flux-corrected transport (mFCT) code which incorporates a background magnetic field and anisotropic pressure. The anisotropic pressure is modeled with a polybaric pressure model (Stasiewicz, 2004). We find that the resultant turbulent state of the interstellar matter (ISM), driven purely by time-dependent heating and cooling with polybaric pressure is similar in nature to the one driven by external forcing such as by supernova feedback. This demonstrates the universal nature of the turbulence in ISM, which is largely isobaric and incompressible. The findings of our analysis is consistent with the contemporary status of knowledge about the multiphase nature of the ISM with volume and mass fractions of the various components of the ISM that is warm, cold, and unstable neutral matter (UNM) in the ranges reported by various numerical and observational analysis.Though the strength of the background magnetic field only marginally affects the overall evolution, the ratio of the parallel and perpendicular pressures can considerably alter the mass and volume fractions of the three phases, which can affect the overall evolution of the TI in the long run.
Magnetohydrodynamics(1964) — Interstellar medium(847) — Interstellar phases(850) —Interstellar dynamics(839)
1 Introduction
Thermal instability (TI) is a physical process responsible for non-gravitational condensations in various plasma environments such as solar corona, solar wind, magnetosheath, interstellar medium (ISM), intracluster medium (ICM) etc. When thermal equilibrium becomes unstable, it can change the structure and dynamics of the medium at a small scale resulting in condensations of high density and low temperature such as clouds and filaments, known as the precursors to star formation. First introduced by Parker (1953) and then by Field in his seminal work (1965), TI is known to explain the existence of multiphase structure of the ISM and mass flow between different intermediate structures, which has very significant effects on the rate of star formation (Chieze, 1987).
The multiphase model of the ISM is a well-studied physical model, which was thought to be consisting of two stable phases – warm and cold, formed out of neutral atomic hydrogen (Field et al., 1969). Studies have later confirmed the existence of these two phases – diffuse warm neutral medium (WNM) and dense cold neutral medium (CNM), believed to be maintained by two thermodynamical states of neutral hydrogen (Wolfire et al., 1995, 2003). Based on earlier studies, it has been argued that large amount of thermally unstable matter is not allowed to exist in theoretical models of global ISM (Kalberla & Haud, 2018) and even if exists, it has to be a transition phase. However, it has been subsequently found, in numerical simulations (Kritsuk et al., 2017; Kim et al., 2023) as well as observationally (Kalberla & Haud, 2018; Murray et al., 2018; Heiles & Troland, 2003; Roy et al., 2013) that a large fraction of the ISM is actually in the unstable neutral media (UNM), comprising about $\sim 20-40\%$, both in volume and mass fractions, making the ISM actually a three-phase medium. As such, we do not have sufficient physical understanding of the UNM phase, the presence which definitely affects the status of both WNM and CNM phases. We should note that besides these three phases, there is also a hot ionized phase, known as the hot ionized medium (HIM), existing at a much higher temperature, believed to be maintained by supernovae heating and a warm ionized phase, maintained by hot B-type stars, known as warm ionized medium (WIM) (McKee & Ostriker, 1977). We in this work, however, do not consider the WIM or HIM phase in our analysis.
Starting from Field’s work (1965), there have been various one-dimensional (1D) studies of TI in both linear and nonlinear regimes with different background conditions. Hennebelle and others (1999; 2000) have observed that transonic converging flows of warm gas can produce long-lived (more than cooling time) condensations of cold gas and in the presence of a magnetic field, in which condensation forms along the field direction due to realignment of fluid flow if the angle between original flow and field is small enough. Another concurrent study of shock propagation into WNM and CNM observed fragmentation of thin dense, thermally collapsed layer into molecular clouds (Koyama & Inutsuka, 2000). They proceeded with several further simulations of TI in one, two, and three dimensions (Koyama & Inutsuka, 2000, 2002; Inutsuka, 2005). There have been many other studies which investigated propagation of a shock front in a thermally unstable gas driving two-phase turbulence in the ISM (Koyama & Inutsuka, 2002; Audit & Hennebelle, 2005), effect of kinematic viscosity on nonlinear turbulent fluid flow (Inutsuka, 2005; Koyama & Inutsuka, 2006), and nonlinear evolution of TI for converging flows of WNM in presence of a magnetic field (Inoue & Inutsuka, 2008, 2009). Koyama & Inutsuka (2004) included isotropic thermal conduction term and put forward the so-called ‘Field condition’ which limits the simulation cell size to avoid artificial fragmentation of the media. In a magnetized background, both pressure and thermal conduction become anisotropic, which leads to the formation of long filamentary condensations along the field direction (Choi & Stone, 2012). Recently, Jennings & Li (2021) have investigated the TI for multiphase gas in the ISM with a Gaussian random field (GRF) density perturbation, in the presence of realistic conduction. Apart from the ISM, there are several other studies on TI in ICM (Parrish et al., 2009; Sharma et al., 2010; McCourt et al., 2011) and solar corona regions (Soler et al., 2011; Antolin et al., 2022; Claes & Keppens, 2019). Falle et al. (2020) have reinvestigated Field’s linear theory of TI using modern wave interaction and stability analysis, providing a clearer representation of the underlying physical mechanisms. That being so, the study of TI with different initial conditions and new physical processes remains always an interesting area for understanding the balance of cooling and heating in the formation of multiphase structures in various astrophysical environments.
We, in this work, perform an inviscid two-dimensional (2D) simulation of the TI in the ISM, with a background magnetic field with pressure anisotropy. We carry out our simulations through a multi-fluid, 2D flux-corrected transport code (mFCT), which takes care of any possible numerical diffusion. We use a polybaric pressure model in our simulation, where we handle the pressure anisotropy with an anisotropy parameter $a_{p}$ (Stasiewicz, 2004, 2005, 2007). With the polybaric pressure equation, we can model the ISM pressure as polytropic in nature, which is supported by observations (Spaans & Silk, 2000) in contrast to the widely used Chew-Goldberger-Low (CGL) double adiabatic pressure model (Chew et al., 1956), which requires very slow plasma flow so that the parallel and perpendicular pressures become completely disconnected, a condition very difficult to be justified (Kulsrud, 1983). The pressure anisotropy parameter is assumed to be constant which is also supported by the fact that variable pressure anisotropy would require the magnetic field to be strongly dependent on density, which however is not supported by observational data for the ISM density range $\sim 0.1-100\,{\rm cm}^{-3}$ (Troland & Heiles, 1986). The polybaric model has been successfully applied to study large amplitude magnetosonic oscillations in Earth’s magnetosheath region. As far as the kinematic viscosity is concerned, we can find many contemporary simulations with inviscid fluid (Padoan et al., 2016; Piontek & Ostriker, 2005; Sharma et al., 2010; Audit & Hennebelle, 2005; Soler et al., 2011; Inoue et al., 2006; Hennebelle & Audit, 2007; Kobayashi et al., 2022, 2020; Inoue & Inutsuka, 2012). We would also like to note that our analysis does not include any star formation activity (Kim et al., 2014, 2023) or supernova feedback term (Kim & Ostriker, 2017) which essentially requires external forcing. Our analysis is to assess the thermal evolution of the ISM and to see whether we can arrive at the similar outcome regarding distributions of WNM, UNM, and CNM with realistic pressure profile and time-dependent heating/cooling, without considering any external forcing terms. We further note that while it is now well-known that supernova feedback and other external forcing terms can drive shock fronts through the ISM; from simulation (Kim et al., 2014) and observational studies (Heiles & Troland, 2003) on $21\,\rm cm$ neutral hydrogen emission, it is observed that the so-called spin temperature and kinetic temperature remain almost equal in the CNM and UNM regions indicating near sonic or subsonic turbulent velocity, ruling out formation of any shock front, which is consistent with our analysis. As we shall see that the nature of the turbulence driven by thermal instability alone is quite similar in behaviour to those driven by external forcing (Kobayashi et al., 2022; Kritsuk et al., 2017; Hennebelle & Audit, 2007; Saury et al., 2014; Kim & Ostriker, 2017) such as converging shocks or supernova feedback, which indicates toward a universal nature of the turbulence, irrespective of being driven by external shock fronts or time-dependent heating and cooling.
The organization of this paper is as follows. In Section 2, we discuss our plasma model along with description of the polybaric pressure equations. We also discuss in this section, why kinematic viscosity may not be too much relevant in the parameter regime that we have studied. In Section 3, we describe our numerical model and briefly also explain the numerical strategy that we have adopted. Section 4 is devoted to the analysis of the linear phase of the TI, carried out with a scaled-down version of our code in 1D, the findings of which also serve as benchmarking results for our code. The primary results of our analysis i.e. the results of fully nonlinear 2D polybaric simulation of TI are presented in Section 5, where we also present our important observations and findings in various subsections. Finally, in Section 6, we conclude.
2 Ideal MHD equations with radiative heat-loss
The conservative form of the ideal MHD equations, which can be written in general as
$\frac{\partial{\cal F}}{\partial t}+\nabla\cdot\mathbb{F}=0,$ (1)
where
${\cal F}=(\rho,\rho\bm{v},{\cal E},\bm{B})^{\prime}$ (2)
is a column vector whose components are mass density $\rho$, momentum density $\rho\bm{v}$, total energy density ${\cal E}$, and magnetic field $\bm{B}$. The quantity $\mathbb{F}$ represents the corresponding flux. The individual equations making up the above model are the continuity, momentum, energy equations, and Faraday’s law combined with Ohm’s law. The compressible fluid equations are given by
$\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\bm{v})$ $=$ $0,$ (3)
$\rho\frac{d\bm{v}}{dt}$ $=$ $\bm{J}\times\bm{B}-\nabla\cdot\mathbb{P},$ (4)
$\frac{\partial{\cal E}}{\partial t}+\nabla\cdot\mathbb{F}_{\textrm{en}}$ $=$ $-\rho\mathfrak{L},$ (5)
$\frac{\partial\bm{B}}{\partial t}+\nabla\cdot\mathbb{F}_{B}$ $=$ $0,$ (6)
where $d/dt\equiv\partial/\partial t+\bm{v}\cdot\nabla$, $\mathbb{P}$ is the anisotropic pressure tensor, and
$\bm{J}=\frac{1}{\mu_{0}}\nabla\times\bm{B}.$ (7)
The lone term on the right hand side of Eq.(5) is the radiative heat-loss term with $\mathfrak{L}$ as the total heat-loss function (cooling and heating) (Field, 1965). The total energy per unit volume ${\cal E}$ is given by
${\cal E}=\frac{1}{2}\rho v^{2}+\frac{p}{\gamma-1}+\frac{B^{2}}{2\mu_{0}},$ (8)
where the terms on the right represent respectively the kinetic energy, internal (thermal) energy, and magnetic energy. The corresponding energy flux $\mathbb{F}_{\textrm{en}}$ can be written as
$\mathbb{F}_{\textrm{en}}=\left(\frac{1}{2}\rho v^{2}+\frac{\gamma}{\gamma-1}p+\frac{B^{2}}{\mu_{0}}\right)\bm{v}-\frac{1}{\mu_{0}}\bm{B}(\bm{B}\cdot\bm{v})+\bm{Q},$ (9)
with $\bm{Q}$ denoting thermal conductivity. In writing the energy equation, we have expressed the thermal pressure as the average scalar pressure
$p=\frac{2p_{\perp}+p_{\parallel}}{3},$ (10)
where $p_{\perp,\parallel}$ are the parts of anisotropic pressure tensor
$\mathbb{P}=p_{\perp}\delta_{ii}+(p_{\parallel}-p_{\perp})\hat{b}_{i}\hat{b}_{j}.$ (11)
The temperature $T$ is related to pressure through the ideal gas law
$T=\mu\frac{p}{R\rho},$ (12)
[…]
$\mathbb{F}_{B}=\bm{v}\bm{B}-\bm{B}\bm{v}.$ (13)
2.1 Generalized continuity equation form
The flux-conservative form of the above MHD equations can be expressed in a form of generalized continuity equation with a source term, which will actually be used in our computation. This form can be written as
$\frac{\partial{\cal F}}{\partial t}=-\nabla\cdot({\cal F}\bm{v})+S,$ (14)
where we split up the total flux term into the flux of the quantity ${\cal F}$ and a corresponding source $S$. The column vector ${\cal F}$
${\cal F}=(\rho,\rho\bm{v},{\cal E},\bm{B})^{\prime}$ (15)
as before and the source terms are given by
$S$ $=$ $\left(0,S_{\textrm{mom}},S_{\textrm{en}},S_{B}\right)^{\prime},$ (16)
$S_{\textrm{mom}}$ $=$ $-\nabla\cdot\left(\mathbb{P}+\frac{B^{2}}{2\mu_{0}}\bm{I}-\frac{1}{\mu_{0}}\bm{B}\bm{B}\right),$ (17)
$S_{\textrm{en}}$ $=$ $-\nabla\cdot\left[\left(p+\frac{B^{2}}{2\mu_{0}}\right)\bm{v}-\frac{1}{\mu_{0}}\bm{B}(\bm{B}\cdot\bm{v})+\bm{Q}\right]-\rho\mathfrak{L},$ (18)
$S_{B}$ $=$ $-\nabla\cdot(\bm{v}\bm{B}-2\bm{B}\bm{v}).$ (19)
Note the above equations do not have the kinematic viscosity term.
2.2 Radiative heat-loss with thermal conduction
While the radiative heat-loss function $\mathfrak{L}\equiv\mathfrak{L}(\rho,T)$ (Field, 1965) is a function of the local density and temperature, the thermal conductivity term, which can be written, in presence of a magnetic field, as
$\bm{Q}=-K_{\parallel}\nabla_{\parallel}T-K_{\perp}\nabla_{\perp}T,$ (20)
where the directions $\parallel$,$\perp$ , ⟂ refer to that of the equilibrium magnetic field $\bm{B}_{0}$. For interstellar matter (ISM), we follow a model for $K$ and $\mathfrak{L}$, which are prescribed as (Falle et al., 2020)
$K$ $\simeq$ $K_{\parallel}=K_{c}T^{1/2},$ (21)
$\mathfrak{L}$ $=$ $\frac{\rho}{m_{H}^{2}}\Lambda(T)-\frac{1}{m_{H}}\Gamma,$ (22)
where $K_{c}$ is a positive constant and $(\Lambda,\Gamma)$ are respectively radiative cooling and heating functions with $m_{H}$ as the mass of neutral hydrogen atom. These functions can further be expressed as
[…]
$\Gamma$ $=$ $2\times 10^{-26}\,\textrm{erg}/\textrm{s},$ (24)
where $a=10^{7}$ cm3 and $b=1.4\times 10^{-2}$ cm3/K1/2 are positive constants and $f_{1,2}(T)$ are certain functions of $T$,
$f_{1}(T)$ $=$ $-\frac{1.184\times 10^{5}}{T+1000},$ (25)
$f_{2}(T)$ $=$ $-\frac{92}{T},$ (26)
where $T$ is measured in Kelvins. We should note that the equilibrium cooling function is assumed to be always zero
$\mathfrak{L}(\rho_{0},T_{0})\sim 0.$ (27)
2.3 Equation of state
The last piece of the model is to specify the equation of state. The isotropic equation of state can be written as
$\frac{d}{dt}\left(\frac{p}{\rho^{\gamma}}\right)=0.$ (28)
The anisotropic pressure, represented by Eq.(11), needs however two equations of state. At the simplest, the equations of state can be expressed through CGL double adiabatic laws (Chew et al., 1956),
$\frac{d}{dt}\left(\frac{p_{\perp}}{\rho B}\right)$ $=$ $0,$ (29)
$\frac{d}{dt}\left(\frac{p_{\parallel}B^{2}}{\rho^{3}}\right)$ $=$ $0$ (30)
The double adiabatic equations, however, are quite restrictive and can be used theoretically only when the system has very slow variation along the field lines such that the communication having different behaviour at two different points, even along the lines of force, is very little (Kulsrud, 1983). Various theoretical and numerical studies also indicate that to a large extent the equation of state for ISM should be polytropic in nature (Spaans & Silk, 2000), which contradicts the CGL equations of state as in CGL, both parallel and perpendicular pressures have different polytropic indices. To circumvent this, the polybaric equations of state are prescribed as (Stasiewicz, 2004, 2005, 2007)
$p_{\perp}\propto N^{\gamma}B^{\kappa},$ (31)
along with a pressure anisotropy parameter
$a_{p}=p_{\parallel}/p_{\perp}-1,$ (32)
where $\gamma,\kappa$ are certain exponents. This model does not impose any restrictions on adiabacity. In what follows, we shall use the $\gamma=5/3$.
2.4 Inviscid plasma flow
While viscosity is a fundamental characteristic of any fluid including plasma, in many situations the fluid viscosity can be safely neglected which simplifies the underlying computational methods. Many such examples can be found which include studies of ISM, ICM, and intergalactic medium (Padoan et al., 2016; Piontek & Ostriker, 2005; Sharma et al., 2010; Audit & Hennebelle, 2005; Inoue et al., 2006; Hennebelle & Audit, 2007; Kobayashi et al., 2022, 2020; Inoue & Inutsuka, 2012). In any case, the effect of fluid viscosity is negligible in the larger scale, while it begins to become important at smaller scales. For studies involving ISM, such as the present work, the usual parameter of interest for fluid viscosity is the Prandtl number
${\rm Pr}=\frac{\gamma}{\gamma-1}\frac{k_{B}}{m_{H}}\frac{\mu}{K},$ (33)
which is the ratio of the kinetic diffusivity to thermal diffusivity, where $k_{B}$ is the Boltzmann constant and $\mu$ is the coefficient of dynamic fluid viscosity. This number is usually fixed at $2/3$, considering the ISM to be mostly monoatomic gas, which fixes the value for $\mu$ depending on the thermal conductivity $K$. In a turbulent flow, the largest scale at which the effect of fluid viscosity becomes important, is usually the Kolmogorov’s scale (1941) (ideally for incompressible flow)
$l_{{\rm kol}}\sim\left(\frac{\nu^{3}}{\epsilon}\right)^{1/4},$ (34)
where $\nu$ is the kinematic viscosity of fluid given by $\nu=\mu/\rho$ and $\epsilon$ is the average rate of dissipation of kinetic energy per unit mass. As we shall see in Section 5.2.3, the Kolmogorov’s scale in our case can be estimated to be $\sim 0.8\,{\rm pc}$, which comes out to be $\gtrsim l_{{\rm kol}}$ for the parameters we have considered in this work. So, we can safely conclude that the effect of viscosity on the results that are obtained through this simulation work is only minimal.
3 Numerical strategy
3.1 The method of flux-corrected transport (FCT)
In this work, we shall use the method of flux-corrected transport (FCT) to solve our initial value problem of the system of equations given by Eq.(14). The FCT method was first introduced by Boris & Book (1973), which solves the Navier-Stokes equations in two stages – a transport stage and a flux-corrected anti-diffusion stage. We shall however use Zalesak’s scheme (1979) which uses a different limiter to limit the anti-diffusive fluxes. The multidimensional time advancement scheme is implemented through a time-step splitting method known as split-step method, which is a second-order accurate method for small enough time steps.
3.2 Normalization
We now normalize all the physical quantities by their respective equilibrium values (denoted with subscript ‘00’). Specifically, we adopt the following normalization
$\rho/\rho_{0}\to\rho,\quad v/c_{s}\to v,\quad t/t_{0}\to t,\quad l/l_{0}\to l,$ (35)
where $t_{0}$ and $l_{0}$ are the acoustic time and length scales with $c_{s}=l_{0}/t_{0}$. We normalize the pressure as
$p/p_{0}\to p,\quad T/T_{0}\to T,$ (36)
where
$p_{0}=\frac{R}{\mu}\rho_{0}T_{0}=c_{s}^{2}\rho_{0},\quad c_{s}=\sqrt{\frac{RT_{0}}{\mu}}.$ (37)
The magnetic pressure $p_{B}=B^{2}/\mu_{0}$ can similarly be normalised by its equilibrium value $B_{0}^{2}/\mu_{0}$ as the magnetic field is normalised by its equilibrium value $B_{0}$
$\bm{B}/B_{0}\to\bm{B}.$ (38)
The thermal conductivity term is scaled as
$\nabla\cdot\bm{Q}\to-\alpha\nabla\cdot\left(T^{1/2}\nabla T\right),$ (39)
where
$\alpha=\frac{t_{0}T_{0}^{3/2}K_{c}}{p_{0}l_{0}^{2}}$ (40)
is the normalization term and all terms are in normalized form. For interstellar matter (ISM), we use $K_{c}\sim 10^{6}\,{\rm erg/cm/K/s}$ and for equilibrium pressure we have $p_{0}/k=4\times 10^{4}\,{\rm K/cm^{3}}$. For other equilibrium quantities, we take $l_{0}=1\,{\rm pc}$, $t_{0}=1\,{\rm Myr}$, $n_{0}\sim 2\,{\rm cm^{-3}}$, $T_{0}\sim 500\,{\rm K}$, and $\mu=1.27$. With these parameters, we have $\alpha\sim 0.037$.
For $T\sim T_{0}$, a quick calculation shows that $f_{1}\ll f_{2}$, appearing in the cooling functions Eqs.(25,26). So, in the normalized form, we can scale the radiative heat-loss term as
$\rho\mathfrak{L}\to\xi\rho^{2}T^{1/2}f_{2}(T)-\delta\rho,$ (41)
with
$\xi$ $=$ $b\frac{t_{0}T_{0}^{1/2}\rho_{0}^{2}\Gamma}{m_{H}^{2}p_{0}},$ (42)
$\delta$ $=$ $\frac{t_{0}\rho_{0}\Gamma}{m_{H}p_{0}}.$ (43)
For the parameters mentioned above, $\xi\simeq 0.072$ and $\delta\simeq 0.23$, which are the normalization terms for the radiative loss function. Note that when normalized, the function $f_{2}(T)$ becomes
$f_{2}(T)\simeq e^{-0.184/T}.$ (44)
We must now note that as all the quantities are normalized with their respective equilibrium values and at equilibrium, we have $(\rho,T,p,B)=1$. We should however note that at equilibrium
$\mathfrak{L}(\rho_{0},T_{0})\sim 0,$ (45)
which actually fixes some of the parameters, depending on the state of equilibrium that is assumed. In any case, one must have
$\delta=\xi f_{2}(T_{0})$ (46)
in equilibrium.
In this work, we consider two cases – one with zero magnetic field and the other with a constant background $\bm{B}$. Without loss of any generality, we assume that the equilibrium magnetic field lies in the $x$-$y$ plane
$\bm{B}_{0}=\hat{\bm{x}}B_{0}\cos\psi+\hat{\bm{y}}B_{0}\sin\psi,$ (47)
at an angle $\alpha$ to the $x$-direction. In this case, the only effect of $\bm{B}$ is through the anisotropic pressure. We note that with the present notation, for constant $\bm{B}$, we have
$\hat{b}_{i}\hat{b}_{j}=\left(\begin{array}[]{ccc}\cos^{2}\psi&\cos\psi\sin\psi&0\\ \sin\psi\cos\psi&\sin^{2}\psi&0\\ 0&0&0\end{array}\right).$ (48)
We also note that as our simulation domain is constrained in the $x$-$y$ plane, we shall restrict all variations in the $(x,y)$ domain only.
3.3 Numerical cycle
In the next sections, we shall present the results for the linear growth phase of the TI in a 1D case. The nonlinear simulations are run in 2D as mentioned in the previous section. In the 1D case, we use a sinusoidal perturbation, which helps us compare the linear growth phase of the TI with the analytical results (Field, 1965). In the 2D case, we use a small randomized density perturbation state with a perturbation level $\sim 1\%$ of the equilibrium density. The initial state is so chosen that it is slightly unstable to the thermal condensation mode. The use of energy equation is desirable for the purpose of conservation of the total energy. Once we determine the total energy, the $p_{\parallel,\perp}$ and the total pressure $p$ can be calculated using Eqs.(10,32). The temperature then can be determined from the scalar (averaged) pressure by using the normalised relation
$T=\frac{p}{\rho}.$ (49)
We use the FCT method to solve Eqs.(14-19) in a uniform grid of size $(1000\times 1000)$ with periodic boundary conditions.
3.4 The characteristic scales
At this point, it is useful to calculate the characteristic time and length scales of the problem. Based on the linear instability analysis already outlined before (Field, 1965), one can define the critical wave number for TI to grow, which provides a characteristic length scale, known as the so-called Field length (Falle et al., 2020; Koyama & Inutsuka, 2004), which can be defined as
$\lambda_{F}=\sqrt{\frac{KT}{n^{2}\Lambda}},$ (50)
where $K=K_{c}T^{1/2}$. The Field length defines the minimum scale that is required for the TI to grow or else, any density perturbation smaller than the Filed length will be quickly erased by the thermal conductivity. Though Koyama and Inutsuka (2000) have used a simplified version of the heat-loss function (Wolfire et al., 1995) to calculate $\lambda_{F}$, one can also use the heat-loss function as given by Eq.(23). For the parameters used in this work, the Field length comes out to be $\lambda_{F}\sim 0.24-0.35\,{\rm pc}$ for $T\sim 500-1500\,{\rm K}$. As we can see, our computational grid of cell size
$\Delta x=2\pi\times 10/1000\simeq 0.063\,{\rm pc}$
can very effectively resolve the Filed length. We also note that our time step size $\Delta t\lesssim 0.01$ is quite within the CFL limit
$t_{{\rm CFL}}\sim C_{{\rm CFL}}\Delta x/c_{s},\quad C_{{\rm CFL}}<1.$ (51)
Next, we calculate the cooling time scale as (Koyama & Inutsuka, 2004)
$t_{{\rm cool}}\sim\frac{kT}{(\gamma-1)}{\rm min}\left(\frac{1}{\Gamma},\frac{1}{n\Lambda}\right)\simeq 0.16\,{\rm Myr}.$ (52)
This can also be used to define the cooling length of the system as (Koyama & Inutsuka, 2004)
$\lambda_{{\rm cool}}\sim c_{s}t_{{\rm cool}}\sim 0.3\,{\rm pc}.$ (53)
We can also define a conduction time scale as (Koyama & Inutsuka, 2004)
$t_{{\rm cond}}\sim\frac{nk}{(\gamma-1)K}\Delta x^{2}\sim 0.022\,{\rm Myr}.$ (54)
We can now safely conclude that our simulation very comfortably resolves the characteristic time and length scales of the problem.
Finally, we do a consistency check by calculating Jean’s length, which puts the upper limit for any structure to be gravitationally stable. Though we have not considered self-gravity in our model, calculating Jean’s length justifies our a priori assumption that gravitational effect may be safely neglected. For parameters already mentioned before, the Jean’s length can be calculated as
$\lambda_{J}=\sqrt{\frac{\pi c_{s}^{2}}{G\rho}}\sim 220\,{\rm pc}.$ (55)
As we shall see in Section 5, considering the highest density condensation of the order of $\sim 10\rho_{0}$ and temperature drop of about $\sim 0.1T_{0}$ in the CNM, the local Jean’s length in the condensations may drop to about
$\lambda_{J}^{{\rm local}}\sim 26\,{\rm pc},$ (56)
both of which are considerably larger than the average condensation size (see Fig.2), which is consistent with our assumption of negligible gravitational effect in the model.
4 The linear growth phase
We know that the linear thermal instability in the absence of magnetic field can essentially grow in two different ways – the so-called condensation mode and the growing wave mode. As the results are quite well known (Field, 1965), we just mention here the instability conditions without derivation. In the absence of thermal conduction, these instability conditions for condensation and wave modes are respectively the isobaric and isentropic instability conditions, which in the normalized forms can be written as
Figure 1: Linear evolution of condensation and wave modes with their growth rates estimated from simulation. (a) both modes unstable (top left), (b) only condensation mode unstable (right). The corresponding growth rates $\varpi_{c,w}$ are shown in the figure, as determined from the simulation. The blue colored graphs (solid and circle) indicate data from simulation and the red colored solid lines indicate an exponential fitting corresponding to the linear growth phase.
$\mathfrak{L}_{T}-\mathfrak{L}_{\rho}$ $<$ $0,$ (57)
$\mathfrak{L}_{T}+\frac{1}{\gamma-1}\mathfrak{L}_{\rho}$ $<$ $0,$ (58)
where $\mathfrak{L}_{T,\rho}=\partial\mathfrak{L}/\partial(T,\rho)|_{\rho,T}$. With thermal conductivity, these conditions become
$k^{2}K+\mathfrak{L}_{T}-\mathfrak{L}_{\rho}$ $<$ $0,$ (59)
$k^{2}K+\mathfrak{L}_{T}+\frac{1}{\gamma-1}\mathfrak{L}_{\rho}$ $<$ $0$ (60)
Theoretically, one can choose different values for $\mathfrak{L}_{T,\rho},K,k$ to make any or all of the above conditions satisfied. For example, for $\gamma=5/3$, we can easily find the following conditions for various instabilities (Field, 1965).
(a) both $\begin{array}[]{lcll}\textrm{(a) both unstable}&:&\mathfrak{L}_{T}<0,|A|<\mathfrak{L}_{\rho}<{\displaystyle\frac{2}{3}}|A|,&A=k^{2}K+\mathfrak{L}_{T},\\ \textrm{(b) only wave}&:&\mathfrak{L}_{\rho}<0,-B<\mathfrak{L}_{T}<-B-{\displaystyle\frac{5}{2}}\mathfrak{L}_{\rho},&B=k^{2}K-\mathfrak{L}_{\rho},\\ \textrm{(c) only condensation}&:&\mathfrak{L}_{\rho}>A,&A>0,\\ &&\mathfrak{L}_{\rho}>{\displaystyle\frac{2}{3}}|A|,&A<0.\end{array}$ (61)
In Fig.1, we have shown the evolution of a single wave in the 1D case for two cases– both modes unstable and only condensation mode unstable, along with the corresponding numerically estimated growth rates $\varpi_{c,w}$ of the condensation and the wave modes. In all the panels of Fig.1, the blue colored curves (solid and circles) denote the results from simulation and the red colored solid line shows an exponential fitting for the linear growth phase. The left panel of Fig.1 shows the evolution of a single wave when both wave and condensation modes are unstable. While the primary figure shows the evolution of the condensation mode with a growth rate of $\varpi_{c}\sim 3.57\times 10^{-3}$ (normalized), the inset shows the evolution of the wave mode with a growth rate of $\varpi_{w}\sim 1.52\times 10^{-3}$, determined from the actual growth of amplitude of the wave after removal of the condensation growth. The right panel of Fig.1 shows the evolution of the condensation mode with a growth rate of $\varpi_{c}\sim 5.29\times 10^{-3}$. The inset shows a zoomed out portion of the initial evolution, where one can clearly see a stable wave mode with only the condensation mode growing. The evolution of the wave mode, corresponding to (b) in Eq.(61) is found to have a growth rate $\varpi_{w}\sim 1.3\times 10^{-3}$ (not shown in figure). The analytically calculated corresponding growth rates $\left(\varpi_{c,w}^{{\rm anal}}\right)$ for these cases are respectively (a) $\varpi_{c}^{{\rm anal}}\sim 3.56\times 10^{-3}$, $\varpi_{w}^{{\rm anal}}\sim 1.52\times 10^{-3}$, (b) $\varpi_{w}^{{\rm anal}}\sim 1.31\times 10^{-3}$, (c) $\varpi_{c}^{{\rm anal}}\sim 5.27\times 10^{-3}$. As we can see, the growth rates determined from the simulation are almost in complete agreement with the analytically calculated ones.
5 Nonlinear polybaric evolution
In this section, we are going to present our primary simulation results with polybaric pressure under a constant magnetic field. All the simulations are carried out in a 2D geometry with a simulation box size of $(20\pi\times 20\pi)\,{\rm pc}^{2}$. The simulation is started with random initial perturbation in the density map, which is slightly unstable to the thermal condensation mode. The perturbation amplitude is set at $\sim 1\%$ of the equilibrium density.
5.1 No magnetic field
In Fig.2, we have shown the evolution of the density map with scalar pressure from a random initial perturbation, which is unstable to the thermal condensation mode. The plots show the thermal condensation in density resulting in formation of cold neutral matter (CNM). The small rectangular box in each panel indicates the area, which is used to calculate the initial (linear) growth rate of the instability. The linear growth of the TI is shown in the right panel of the bottom row of Fig.2. The blue circles indicate the density evolution (within the rectangular box) and the red curve is an exponential fit corresponding to the linear regime. For the ISM parameters already mentioned before and the equilibrium state that we have used, the maximum linear growth rate (with respect to perturbation wave number) that is possible, comes out to be $\varpi_{{\rm max}}^{{\rm theory}}\sim 1.03\times 10^{-2}\,{\rm Myr^{-1}}$, which we are expected to observe during a simulation with random perturbation. As shown in Fig.2 (right panel, bottom row), the linear growth rate in this case is found to be $\varpi^{{\rm simul}}\sim 1.93\times 10^{-2}\,{\rm Myr^{-1}}$, which closely agrees with the theoretical value. The density $\rho$ then increases from an equilibrium value of unity (normalized to the equilibrium value $\rho_{0}$) to about an average value of $\sim 180$ during the nonlinear phase.
Figure 2: The evolution of the density map under TI. The timestamp in different simulation panels are inscribed within each panel. The initial state of the simulation is a random density perturbation which is unstable to the thermal condensation mode. The small rectangle shown in each of the top four panels indicates a density condensation area, which is used to calculate the growth rate $(\varpi)$, shown in the right panel of the bottom row. The left panel of the bottom row shows temperature perturbations corresponding to density perturbations at $200\,{\rm Myr}$.
As one can see that the nonlinear evolution starts at about $\sim 500\,{\rm Myr}$ from a very small value (just $\sim 1\%$of the equilibrium density). Naturally, the instability will only linearly grow in the initial phase. We should note that if a higher perturbation amplitude is introduced at the initial stage, the nonlinear evolution will occur sooner. As such, the onset of nonlinear evolution and hence the time required to attain the final condensation stage is quite arbitrary and may depend on the particular simulation concerned. What is however invariant is the linear growth rate of the instability $\varpi$, which is quite consistent with the analytical estimate, as already shown in the previous section and in the beginning of this section. We also note that the nonlinear evolution never attains a completely steady state and fluctuates around a mean value and hence the size and structure of the CNMs will fluctuate with time but the overall density structures have a nonlinear evolution timescale of $\sim 10\,{\rm Myr}$. Note that the density shown in this plot is a cumulative density within that rectangle. In left panel of the bottom row of Fig.2, we have also shown the normalized temperature map $(T/T_{0})$ at a timestamp of $200\,{\rm Myr}$, which clearly shows the anti-correlation between the corresponding density map and formation of CNM and warm neutral matter (WNM), which further confirms the process of thermal condensation with pressure equilibration.
5.2 Constant magnetic field
We note that experimental observations suggest that the magnetic field strength is almost independent of density in the range of $0.1-100\,{\rm cm^{-3}}$ or in other words the density depleted regions (warm neutral matter or WNM) and the condensations (cold neutral matter or CNM) have roughly the same field strengths (Troland & Heiles, 1986). We therefore restrict our analysis to a constant background magnetic field $\bm{B}\equiv B\hat{\bm{x}}$ in the $x$-direction. Following Stasiewicz (2004, 2005, 2007), we also introduce a pressure anisotropy parameter
$a_{p}=p_{\parallel}/p_{\perp}-1,$ (62)
so that depending on the value of the parameter $a_{p}$, one can have different values for $p_{\parallel,\perp}$.
5.2.1 Plasma $\beta$
As the extent of different plasmas of astrophysical interest is quite large, the plasma $\beta$,
$\beta=\frac{p}{B^{2}/(2\mu_{0})},$ (63)
which is the ratio of the plasma pressure to the magnetic pressure, also can have an extremely wide range of values. While the interstellar matter (ISM) has an extremely wide range of $\beta\sim 9\times 10^{-8}-7\times 10^{8}$ (Haverkorn & Spangler, 2013; Ponnada et al., 2022), the intracluster medium (ICM) as such have range $\beta\sim 0.1-10^{7}$ (Kunz et al., 2010; Berlok et al., 2021). The magnetosheath region has a $\beta\sim 0.1-10^{5}$ (Stasiewicz, 2004; Pang et al., 2022; Bandyopadhyay et al., 2022), while the solar corona and solar wind have a typical range of $\sim 10^{-4}-10^{3}$ (Gary, 2001; Rodríguez Gómez et al., 2019; Huang et al., 2023). We have therefore restricted our study to values of $\beta=10^{4}$ and $5$, which we refer to as the ‘weak’ and the ‘strong’ field.
Figure 3: Density map for $a_{p}=0$ and $\beta=10^{4}$ (‘weak’ field) for different phases of the TI (top four panels). While the top three panels show the evolution for $a_{p}=0$, the right panel of the second row shows the same for $a_{p}=3$. The bottom row shows the evolution at the end of $850\,{\rm Myr}$ for $a_{p}=0$ and $3$ for the case of ‘strong’ field $(\beta=5)$.
   Refer to caption          Refer to caption
Figure 4: The velocity field for $\bm{B}=0$ (top row) and ‘strong’ field (second row) with $\beta=5$ and $a_{p}=3$ at different time stamps. The local Mach number of the plasma flow for the case $\beta=5$ at $850\,{\rm Myr}$ is shown in the left panel of the third row. The bottom two panels on the right show the power spectrum density (PSD) of the velocity field for $\bm{B}>0\,(\beta=5)$ (top) and $\bm{B}=0$ (bottom). The respective scalings with the spatial wave number are inscribed in the panels. The left bottom panel shows the compensated PSDs of the two cases. The yellow and black colored spectra in both panels indicate the solenoidal and compressible parts that comprise the total spectra.
5.2.2 Thermal condensation
In Fig.3, we have shown the evolution of the density map for the ‘weak’ field $(\beta=10^{4})$ for the parameter $a_{p}=0$ and $3$ (top four panels). The initial state is the same as before (as in the case of scalar pressure). What we can see from the figure is that the initial evolution of density is predominantly along the ambient magnetic field and the final condensations are also developed parallel to the magnetic field. However, as the evolution becomes fully nonlinear, the condensations become diffused compared to the similar map for scalar pressure (see Fig.2) though the final condensations seem to to be aligned with the ambient magnetic field. While all the top four panels in Fig.3 refer to the ‘weak’ field case for $a_{p}=0$ and $3$, the lowermost panel of Fig.3 shows the evolution of density after $850\,{\rm Myr}$ for $a_{p}=0$ and $3$ for the ‘strong’ field case. We can see that for comparatively higher $p_{\parallel}$ (larger $a_{p}$), we have slightly lower condensations.
We now analyse what exactly happens in the case of a strong field. In particular, we choose the case for $\beta=5$ and $a_{p}=3$. We have chosen to plot the velocity field vectors in Fig.4 which is superimposed on a velocity field map to highlight the plasma flow. Note that in Fig.4, the lengths of the velocity vectors are kept uniform for clarity, while the background velocity field indicates the strength of the net plasma flow at that point. The top row of Fig.4 shows the velocity field vectors and map for scalar pressure when there is no ambient magnetic field. We can see that in this case, the plasma velocity is more or less ordered, driven by the density condensations. During the fully developed nonlinear stage of the evolution at $t\sim 850\,{\rm Myr}$, the bulk plasma flow is quite ordered except at the site of the condensations. As we can see that the density condensations are also closely followed by higher plasma flow into the regions of higher mass densities. In contrast to this, when there is a magnetic field present, initially most of the plasma flow is along the magnetic field lines. The low density areas are marked by small-magnitude perpendicular plasma flow, as expected. This also causes initial density striations to form along the magnetic field lines. We note that the immediate effect of a magnetic field is to restrict the plasma flow in the transverse direction, which points to the fact that condensation always occurs along the magnetic field lines as the parallel velocity is mostly unrestricted. However, even a small transverse velocity component can make a considerable difference in the final condensation state (Hennebelle & Pérault, 1999). As a result, we can expect that the thermal instability will be stifled due to the flow restrictions imposed by the magnetic field. However, as the total energy content must be conserved, we certainly expect a field aligned, more diffused condensation state and the resultant velocity field will not be highly ordered leading to a turbulent-like plasma flow, as can be seen from the right figure of the second row of Fig.4 (see next subsection).
5.2.3 Velocity fluctuation and isobaric turbulence
In absence of any induced (external) large-scale velocity fluctuations, say from a supernova feedback term, the velocity field is developed self-consistently due to the formation of CNM and WNM. Besides, as our system is dissipation-less, we do not expect to see any shock front and the velocity is supposed to stay subsonic with respect to the local sound velocity at any point. In Fig.4, we show the local Mach number $M=v/\sqrt{T}$ (left panel, third row), the ratio of the net velocity to that of the local sound speed at a given point within the computational domain. As we can see, the local Mach number always stays well below $1.0$ in the entire computational domain and the plasma flow remains subsonic. We can compare this to the so-called turbulent Mach number $M_{t}$, determined through observation and defined as
$M_{t}=\left[4.2\left(\frac{T_{k,\rm max}}{T_{s}}-1\right)\right]^{1/2},$ (64)
where $T_{k,\rm max}$ is maximum kinetic temperature and $T_{s}$ is the spin temperature which is basically the HI excitation temperature. Recent simulation studies (Kim et al., 2014) as well as in observational analysis (Heiles & Troland, 2003) show that for temperature ranges relevant for the CNM and WNM, the ratio $T_{k,\rm max}/T_{s}\gtrsim 1$, which indicates a near sonic or subsonic turbulent velocity, which is consistent with our findings. It should also be noted that a violent initial perturbation can push this Mach number close to unity making a shock-like interface between the CNM and WNM.
We note that the basic nature of thermal condensation mode is isobaric and the resultant nonlinear state should also show the nature of isobaric turbulence. We analyse the nature of these turbulence through a spatial power spectrum density (PSD) of the velocity field for both zero and strong magnetic field. In Fig.4, we show this spatial PSD for the case $\bm{B}>0,\beta=5$ (third row, right panel) and for $\bm{B}=0$ (bottom row, right panel) with respect to the spatial frequency (expressed as the inverse of the pixel of the velocity field image). What is shown is the 1D power spectrum density of the velocity field calculated using the Turbustat package (Koch et al., 2019). Note that for the 1D PSD, we have $P(k)\equiv E(k)\propto k^{-\kappa}$, where various values of the index $\kappa$ determine the nature of the turbulence. In the figure, the velocity PSDs are shown in the right panels of the third and fourth rows, respectively for the case when $\bm{B}>0$ (with $a_{p}=3$) and $\bm{B}=0$. In both these panels, the blue and red colored spectra (with relative error region marked) show the 1D PSD with a suggestive fitting (thick, solid, black line). The yellow colored spectra in both panels show the solenoidal part of the spectrum and the black colored part shows the compressional part (see the next paragraph for explanation). The left panel of the bottom row shows the compensated spectra for both cases (blue for $\bm{B}>0$ and red for $\bm{B}=0$) for more clarity.
A few points should be noted while interpreting the power spectra shown in Fig.4. First, as our system is dissipation-less, the entire range of the power spectra is inertial as the so-called viscous dissipation range is practically zero. Which is why also, the spectra do not develop any marked transitions indicating the dissipation range. The spectra however show two distinct scalings, one which is marked with black solid line and the other with a green solid line in the figures. For the range shown in the figure by two vertical dashed lines, it is interesting to note that while the scalar pressure (zero $\bm{B}$) velocity field shows a scaling more akin to that of Burger with $E(k)\propto k^{-2}$, with a strong magnetic field $(\beta=5)$, the turbulence scales more like the classical incompressible Kolmogorov scaling with $E(k)\propto k^{-1.69}$. So, the effect of a magnetic field is to make the TI-induced isobaric turbulence more like an incompressible turbulence, at least at larger scales. It is also worthwhile to note that recent observations by Voyager I, which indicate incompressible turbulence in the very local interstellar medium (VLISM) (Zhao et al., 2020). To further confirm our findings, we have decomposed the PSD of the velocity field into two parts namely compressional and solenoidal parts through a Helmholtz-Hodge decomposition (Bhatia et al., 2013)
$\bm{v}=-\nabla\phi+\nabla\times A,$ (65)
where the $-\nabla\phi$ is the compressional part of the velocity field and the other solenoidal. From the figure, it is now amply clear that the total velocity PSDs closely follow their corresponding solenoidal parts for the most part which suggests that the flow is largely incompressible. However, there is marked difference between the two cases when $\bm{B}>0$ and $\bm{B}=0$. If we denote the deviation of the total PSD from the respective solenoidal PSD by $\Delta U_{k}$ for any $k$, we find that the mean deviation in case of scalar pressure $(\bm{B}=0)$ is twice as large as the same for the polybaric pressure $(\bm{B}>0)$ with $a_{p}=3$
$\left(\overline{\Delta U_{k}}\right)_{{\rm scalar}}\gtrsim 2\left(\overline{\Delta U_{k}}\right)_{{\rm polybaric}},$ (66)
which indicates that with magnetic field the flow becomes more incompressible, closely resembling Kolmogorov scaling. As mentioned, the turbulence remains largely isobaric in nature, which is confirmed by the pressure distribution function, shown in the bottom panel of Fig.5. The $\delta$ function-like behaviour of the pressure distribution function indicates that pressure remains isobaric. This turbulence, driven only by time-dependent heating and cooling, is surprisingly similar in nature to the turbulence driven by external forcing such as by supernova feedback or similar activities, creating a shock front, also known as the evaporation front (Kobayashi et al., 2022) in the WNM-CNM interface. Our analysis confirms this universal nature of turbulence in the ISM.
As already mentioned in Section 2.4, our flow is inviscid. In that sense, the entire spectrum range in Fig.4 can be assumed to be inertial from which we can estimate an equivalent of Kolmogorov’s scale to be $l_{{\rm equiv}}\sim 0.72\,{\rm pc}$ corresponding to a maximum wavenumber $k\sim 0.55\,{\rm pix}^{-1}$. For a Prandtl number ${\rm Pr}=2/3$, the dynamic viscosity is about $\mu\sim 0.072$ for the parameters that we have considered in this work. The average rate of dissipation of kinetic energy per unit mass $\epsilon$ can be estimated to be
$\epsilon\sim\frac{\left(\overline{u_{{\rm rms}}}\right)^{3}}{l_{{\rm scale}}}\sim 5.5\times 10^{-7}\,{\rm erg/gm/s},$ (67)
where $u_{{\rm rms}}$ is the r.m.s velocity of the plasma flow and $l_{{\rm scale}}\sim 0.063\,{\rm pc}$ is the smallest scale that is resolved in our simulation. With this, the Kolmogorov’s scale [Eq.(34)] is estimated to be
$l_{{\rm kol}}\sim 0.66\,{\rm pc},$ (68)
which is $<l_{{\rm equiv}}$, the smallest structure that is generated in our simulation. So, we can see that viscosity plays only a minimal role in determining the overall structure formation due to TI.
It has been stated earlier that the presence of these two different scalings of the velocity PSD may be due to numerical diffusion (Kobayashi et al., 2022; Hennebelle & Audit, 2007). But, our analysis shows (as our code is specifically designed to take care of this very numerical diffusion, which is also the characteristics of our FCT code), we can safely argue that the presence of two scalings is not due to numerical diffusion, rather it indicates transition of the turbulence behaviour from either Kolmogorov’s or Burger’s at smaller scales.
5.2.4 Probability density function (PDF)
In Fig.5 (top row), we have shown the probability density function (PDF) for the scalar pressure and anisotropic pressure with $\beta=5$ and $10^{4}$ and $a_{p}=0$ and $3$ at $850\,{\rm Myr}$. As we can see, without any ambient magnetic field, density distribution shows clear indication of two distinctly separated phases with a modest amount of unstable gas in the intermediate region. However, with a background magnetic field, the growth of condensation is restricted, leading to a wide dispersion in the lower density regions. While for ‘weak’ magnetic field the condensations are roughly same for both $a_{p}=0$ and $3$, for the ‘strong’ magnetic field, we see a marked difference between the final condensation distribution; specifically, when $p_{\parallel}$ is considerably higher than $p_{\perp}$, the thermal condensation is slightly suppressed, resulting in a greater proportion of unstable gas compared to equal the case when $p_{\parallel}\sim p_{\perp}$. The middle panel of Fig.5 shows the PDF of thermal pressure (left), from which we can see that pressure-redistribution is quite affected by the magnetic field making the turbulence even more isobaric.
Refer to captionRefer to caption
 Refer to caption   Refer to caption
[…]
In contrast to the above, the mass fraction in the WNM phase stays almost constant at about $\sim 16-17\%$ for both scalar and polybaric pressure, while that in the CNM phase decreases from about $\sim 67\%$ for scalar pressure to about $\sim 42\%$ for $a_{p}=3$ with $\beta=5$. The mass fraction in the UNM phase, however, increases by about $2.5$ times from $\sim 17\%$ for scalar pressure to about $\sim 41\%$ for $a_{p}=3$ with $\beta=5$. We note that all these findings are quite consistent with contemporary observational data, especially for the UNM phase (Heiles_2003; Kalberla & Haud, 2018; Murray et al., 2018). So, the presence of a background magnetic field considerably affects the mass and volume distribution of the WNM and CNM phases, with the magnetic field making the condensations more diffused. Many previous studies have neglected analysis of the long-lasting UNM phase, considering it to be a transient phase formed between the WNM and CNM phases. However, our analysis shows that UNM occupies a very significant amount of fractions both volume and mass-wise and turbulence remains crucial in maintaining this substantial quantity of thermally unstable gas. Similar natures of filling factors were also reported by earlier simulations (Kim et al., 2023; Kritsuk et al., 2017; Audit & Hennebelle, 2005).
[…]
Our analysis of the three ISM phases further confirms the non-transient nature of the UNM and the volume and mass fractions of the multiphase ISM, calculated from our simulation is consistent to the contemporary results (Kim et al., 2023; Kritsuk et al., 2017; Audit & Hennebelle, 2005), which indicates to the fact that with a polybaric pressure profile, even without external forcing, we arrive at a pretty much same results, so far as the structure of the multiphase ISM is concerned. We also note that the background magnetic field makes the size of the WNM (in terms of volume) less than the case with scalar pressure, at the cost of the UNM, leaving the CNM size almost unchanged.
6 Summary and conclusion
This study examines the nonlinear development of isobaric TI due to time-dependent radiative heat-loss in a collisionless warm anisotropic and inviscid plasma using a 2D FCT simulation, for a polybaric equation of state across three different scenarios: unmagnetized, weakly magnetized, and strongly magnetized. The FCT simulation ensures that the simulation is not affected by numerical diffusion (Boris & Book, 1973). The simulation is started with a very small random perturbation in the density map, with an equilibrium, which is slightly unstable to the thermal condensation mode. As time progresses, the TI starts to build up with the linear phase, before entering a quasilinear and finally a nonlinear saturation regime. As such, we establish the linear instability conditions for various modes along with their corresponding growth rates (Field, 1965), which also serves as the benchmarking results for the nonlinear simulation. The effect of pressure anisotropy in the presence of an ambient magnetic field is handled with the polybaric equation of state (Stasiewicz, 2004), helping us to maintain the polytropic nature of the thermal pressure. We also show that for the parameter regime considered in this work, kinematic viscosity can be safely neglected. Apparently, our simulation with the relevant parameters, is able to resolve all the necessary time and length scales.
The important findings of our study can be summarized as:
  • The nature of the thermal instability-driven turbulence is similar in nature to the one driven by external forcing such as by supernova feedback and similar activities (Kobayashi et al., 2022; Kritsuk et al., 2017; Hennebelle & Audit, 2007; Saury et al., 2014). At the turbulent phase, most of the power in the PSD lies within the solenoidal mode, which confirms the incompressible nature of the flow. However, turbulence in the absence of a magnetic field surprisingly scales more like Burgers turbulence, even though no shock is involved in our simulation. In this case also, most of the power in PSD is solenoidal. Our analysis shows the universal nature of turbulent ISM, irrespective of any external forcing.
  • For the ‘strong’ magnetic field, the turbulence scales like the classical Kolmogorov turbulence, which indicates that turbulence remains largely incompressible. We also note similar kind of low frequency, isobaric, incompressible turbulence in the VLISM, as reported recently by Zhao et al. (2020).
  • At the nonlinear saturation regime of the TI, our analysis also shows the ISM to be consisting of three-phase matter, namely WNM, CNM, and UNM. The volume and mass fractions of the three ISM phases are consistent with the contemporary studies which include external forcing and star formation activities (Kim et al., 2023; Kritsuk et al., 2017).
[…]
  • As far as the effect of polybaric pressure anisotropy is concerned, we have observed that the growth of TI is suppressed when the parallel component of the pressure becomes larger than the perpendicular component, leading to a comparatively larger fraction of the UNM. This is indicative of the fact that the anisotropy parameter $a_{p}$ plays a key role in the evolution of the ISM. Obviously, the effect of pressure anisotropy increases with the increase in magnetic field strength. For both ‘weak’ and ‘strong’ fields, the higher is the value of $a_{p}$, the lower is the mass and volume fraction of the condensation. However, the chosen strengths of the magnetic field do not show any pronounced difference in the evolution. We would also like to note that a stronger field is expected to cause defragmentation of the field aligned strands, which may be expected at a much lower $\beta$ for our case (Choi & Stone, 2012).
As a concluding remark, we would like to emphasize that, to our knowledge, this study is the first-ever comprehensive numerical study of TI in the ISM using a flux-corrected transport (FCT) method and also the first study to use polybaric pressure model in investigating TI in the ISM.
One of the authors, HS would like to thank CSIR-HRDG, New Delhi, India for the Senior Research Fellowship (File No: 09/059(0074)/2021-EMR-I)